OpenStudy (aroub):

Dana works in a restaurant. It takes her 4 hours to clean a load of dishes. Her colleague Tim can do the same job in 6 hours. If they work together, what fraction of the load can they clean in one hour?

OpenStudy (aroub):

-Work problem..

OpenStudy (aroub):

5/12 right?

OpenStudy (hoblos):

im one hour dana cleans 1/4 load tim cleans 1/6 load together they clean 1/4 + 1/6 = 5/12 load

OpenStudy (hoblos):

yup it's right

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

and to anticipate the next question, how long will it take them both working together to complete the load? \[\frac{4\times 6}{4+6}=\frac{24}{10}=\frac{12}{5}=2.4\] the reciprocal of the answer to your previous question.

OpenStudy (anonymous):

I like Dana's rate + Tim's rate = Combined Rate ...

OpenStudy (saifoo.khan):

\[\frac 14 + \frac 16 = \frac 1x\]

OpenStudy (anonymous):

5/12